OpenStudy (anonymous):

Please help! I'm unsure how to calculate this: Going against the current, a boat takes 6 hours to make a 120-mile trip. When the boat travels with the current on the return trip, it takes 5 hours. What is the rate of the boat in still water? I know that r = 120/t, but i'm unsure about how to calculate. Here's what I've tried so far:

OpenStudy (anonymous):

it's in still water, so there is no presence of a current. i canceled them out: 6-5 = 1. But, I KNOW that the rate is not 120 mph!

OpenStudy (anonymous):

so then i tried putting the two equations together, so... 5x + 5y = 120 and 6x -6y = 120 11x - y = 240 but obv that doesn't help =/

OpenStudy (anonymous):

there's some other stuff i wrote but it didn't help and the math doesn't "add up" lol>pun

OpenStudy (anonymous):

maybe i need to use a diff approach to the problem?

OpenStudy (anonymous):

d=120 ... d=rt ... 120=(6)*(r+c) ... 120=5*(r-c) ... 6r+6c= 5r-5c ... minus 5r to both sides, and minus 6c to both sides. r=-11c ... 120=6(-11c+c) ... 120=-66c+6c c=-2 r=22

OpenStudy (anonymous):

ok thank you =)